# Evaluating Decision Rules Across Many Weak Experiments

### v1

@@ 4.1 Setup
Importantly, we do not assume that the covariance matrix $\Omega$ is diagonal. In general, $Y$ and $S$ will be correlated at the unit-level outside of any particular experiment (Cunningham and Kim, 2020). For example, highly engaged users ($S$) are also likely to be more retentive ($Y$), regardless of whether interventions that increase engagement also increase retention. Thus, we distinguish between the true associations between the underlying treatment effects and the observed associations between treatment effects estimated from finite experiments (Bibaut et al., 2024). This finite-sample noise amplifies the winner’s curse for proxy metrics and motivates the use of our CV estimator.
@@
$R_{iD_{S}(\mathbf{O}^{i})}=\begin{cases}\tau_{Y}^{i}&\text{if $\hat{\tau}_{S}^{i}>0$}\\ 0&\text{otherwise}.\end{cases}$ (7)
@@
4.2 Closed-Form Analysis
@@
$(\tau_{R}^{i},\tau_{S}^{i})\sim\mathcal{N}\left(\mathbf{0},\Lambda\right).$ (8)
@@
$\mathbb{E}[R_{iD_{S}}]\propto\frac{\Lambda_{12}}{\sqrt{\Lambda_{22}+2\Omega_{22}/M}},$ (9)
which is clearly increasing in the covariance in true treatment effects. [^5] See Appendix 7.1 for derivations of Equations 9-11.
@@
$\mathbb{E}[\hat{R}^{naive}_{iD_{S}}]\propto\frac{\Lambda_{12}+2\Omega_{12}/M}{\sqrt{\Lambda_{22}+2\Omega_{22}/M}}.$ (10)
Therefore, the naive estimator can prefer proxy metrics whose correlation with the north star is driven more by noise than by true treatment effects. A classic example is clickbait: although highly engaged users may be more retentive outside of any particular experiment ($\Omega_{12}$ large), interventions that merely increase clickbait will not have the desired effect on retention ($\Lambda_{12}$ small).
Clearly, the risk of bias is greater when the scale of true treatment effects is very small relative to the unit-level sampling variance, as is often the case in digital experiments (Larsen et al., 2024). This suggests that the naive estimator will skew toward highly sensitive, “lower funnel” metrics such as page visits, as compared to lower funnel metrics that are harder to move (e.g. signups or checkouts) but more strongly correlated with north star business metrics like revenue.
@@
$\mathbb{E}[\hat{R}^{CV}_{iD_{S}}]\propto\frac{\Lambda_{12}}{\sqrt{\Lambda_{22}+2\Omega_{22}/M_{p}}},$ (11)
where $M_{p}=M(P-1)/P$. Like the true reward, this expectation is also insensitive to the noise covariance $\Omega_{12}$ and only depends on the covariance between the true treatment effects. It is also negatively biased due to the division in (11) by $M_{p}<M$, although again this can be mitigated by increasing $P$.
Figure 1: Expected Values of True Reward and Estimators. While the true expected reward only depends on the covariance of true treatment effects, the naive estimator is sloped and can prefer “Bad” proxies to “Good” proxies. The CV estimator is negatively biased (diluted) but ranks proxies correctly.
Figure 1 illustrates these dynamics. In the top panel, we plot the true reward $\mathbb{E}[R_{iD_{S}}]$ as the correlation between true treatment effects ($y$-axis) and measurement error ($x$-axis) vary from -1 to 1.666The other parameter values are set to the same values as in our simulation below. Regions of the same shade represent level sets (i.e., the same value of the reward). As the panel shows, the true expected reward is only sensitive to the correlation in true treatment effects.
@@
Lastly, the bottom panel of Figure 1 plots the expected value of our CV estimator with just two folds. The flat level sets demonstrate that the CV estimator will rank correctly based on the true reward, although their shade is diluted due to the secondary source of bias in the CV reward estimates. In practice, this bias is easily mitigated by increasing the number of folds.
4.3 Simulation Study
@@
As can be seen from (9) and (10), the relative bias of the naive estimator is proportional to the ratio of the covariance in measurement error to the covariance in true treatment effects $\Omega_{12}/\Lambda_{12}$. Letting $\sigma_{S}$ denote the standard deviation of $S$ and $\sigma_{\tau_{S}}$ denote the standard deviation of $\tau_{S}$, the relative bias is therefore proportional to $\sigma_{S}/\sigma_{\tau_{S}}$, the noise-to-signal ratio of $S$. Therefore, as treatment effects on the proxy metric become infinitely small relative to their measurement error, the relative bias of the naive estimator also tends to infinity. This is illustrated in Figure 2, where we simulate the bias of the naive estimator as the scale of the noise $\sigma_{S}$ tends to infinity with $\sigma_{\tau_{S}}$ fixed. As the plot shows, the relative bias of the naive estimator is also unbounded. In contrast, because the CV estimator is biased towards zero, its relative bias is negative for positive rewards and also bounded between $-100\%$ and $100\%$.
Figure 2: Bias of Naive and CV Estimators vs. Noise-to-Signal Ratio of $S$. As the scale of the measurement error in $S$ tends to infinity for a fixed scale of treatment effects on $S$, the bias of the naive estimator tends to infinity, whereas the relative bias of the CV estimator is bounded.
@@
Figure 3: Asymptotics as $M\to\infty$. While the bias of both estimators goes to zero as the number of units per treatment arm $M\to\infty$, the naive estimator is much more biased at small values of $M$, which can lead to the wrong choice of proxy.
@@
Figure 4: Asymptotics as $N\to\infty$. Increasing the number of experiments $N$ without increasing the size of each experiment only increases the bias of the naive estimator.
5 Real-World Application
In this section, we describe the application of our method to advocate for a new experimentation decision rule at Netflix. This work is part of a larger project to develop proxy metrics via meta-analysis, initially described in Bibaut et al. (2024). In that paper, we developed methods for fitting linear structural models of treatment effects using data from past experiments. We deployed these methods to construct an improved proxy metric as a linear combination of component metrics. We then used techniques in this paper to demonstrate the favorability of the resulting metric relative to the previous decision metric. This led directly to the adoption and implementation of the new metric on Netflix’s experimentation platform, where it has been computed for 257 A/B tests as of this writing.
5.1 Setup
The data for this exercise come from 123 past A/B tests at Netflix. We chose these tests based on their similarity to the future tests that we wanted our new proxy metric to inform. Specifically, we focused on tests that aim to improve engagement through in-product interventions, such as improvements to our recommendation algorithms.
Letting $Y$, $Z_{0}$, and $Z_{1}$ denote the north star metric, status quo proxy metric, and the new “challenger” proxy metric, respectively, we estimated the cumulative returns to a handful of candidate launch rules:  [^9] We also experimented with a number of other combinations (e.g. different tie-breakers, requiring statistical significance in both $Z_{0}$ and $Z_{1}$), which we omit for brevity.
  1. 1.
@@ 5.2 Results
Figure 5: Estimated Cumulative Returns to Decision Rules across 123 A/B Tests at Netflix. We applied the methods in this paper to estimate the naive and cross-validated cumulative returns to three different decision rules at Netflix, contrasting two challenger rules to the status quo rule. We find that a new proxy metric would increase cumulative returns by over $33\%$ (p-values between $0.07$ and $0.09$).
@@
Figure 5 shows our results by plotting the naive (“No Folds”) estimate of the cumulative returns (red Xs), alongside our CV estimates under increasingly numerous folds (blue circles). Error bars represent 95% confidence intervals. To maintain confidentiality, we scale all values by baseline gains (the ”No Folds” variant for Rule 1).
@@
As expected, the naive estimator systematically overestimates the cumulative returns of past experiments due to the winner’s curse. Although cross-validation could over-shrink these estimates, the estimates remain stable as we increase the number of folds, indicating that the negative bias of the CV estimator is small in our application.
@@
As the figure shows, our methods can also influence qualitative comparisons between decision rules. The naive ranking places Rule 3 far above Rule 2, while the CV estimator makes the two rules appear comparable. This is likely due to a combination of $Z_{0}$ having a more severe correlated errors problem than $Z_{1}$, as well as the slightly higher Type I error rate incurred by using an either-or launch rule. Altogether, the results demonstrate that the choice of estimator can meaningfully alter both quantitative (i.e. estimates of cumulative impact) and qualitative (the choice of rule) business conclusions.
@@
We find that both “challenger” rules would meaningfully outperform the status quo. In particular, we estimate that adopting Rule 3 would have increased north star metric gains in excess of 33%, with p-values between 0.07 and 0.09 depending on the number of folds. While marginally significant under traditional frequentist thresholds, the large gains at stake and modest implementation costs gave us confidence to recommend Rule 3. These estimated gains, as well as the explainability of our methodology, led directly to the adoption of this new launch rule by decision-makers.
6 Conclusion
@@
This paper proposes to evaluate candidate decision rules via their cumulative returns to business north star metrics in past experiments. In our experience, cumulative returns are an intuitive evaluation metric that resonates with non-technical stakeholders, making them a compelling quantity to report in a broad set of applications. Nevertheless, accurately estimating cumulative returns can be challenging when treatment effects are small relative to measurement error, as is often the case in digital experimentation, including settings like Netflix’s (Bibaut et al., 2024; Larsen et al., 2024; Ejdemyr et al., 2024).
@@
We demonstrate theoretically and empirically that, under such conditions, a cross-validation estimator can estimate the reward of a decision rule effectively. In contrast, naive approaches that reuse the same data for both decision rule selection and evaluation lead to classical winner’s curse biases, even when experimenters are optimizing for proxy metrics rather than north star metrics.
@@
An implicit assumption underlying this work is the independence of treatment effects across tests: The difference in rewards between any two treatment arms within an experiment cannot depend on the launch decisions of other experiments. This assumption could be violated, for example, if experiments implement redundant interventions or if previous launch decisions radically affect the design and impact of future tests. We believe this is a reasonable approximation in our setting, which is characterized by many modest interventions across diverse domains and experiments that are typically planned far in advance. That being said, one valuable direction for future research is to explore estimators that allow for dependence across experiments.
@@
A second, more applied direction in which extend this research is to evaluate other dimensions of experiment decision-making beyond the choice of outcome metric. For example, recent papers have debated whether statistical significance thresholds should be more restrictive or permissive (Kohavi and Chen, 2024; Sudijono et al., 2024). The methodology outlined here provides a principled way for firms to evaluate the returns to different thresholds using actual experiments, and can easily be extended to include costs.
@@
Lastly, another direction for future research is to learn good decision rules rather than pick the best rule from an ex ante-defined set. Here, an interesting trade-off is between optimality and interpretability: a good decision rule should increase the returns to experimentation, but also be coherent enough to be trusted, understood, and remembered by decision-makers with varying statistical aptitudes. We leave this for future work.

### v2

@@ 4.1. Setup
Importantly, we do not assume that the covariance matrix $\Omega$ is diagonal. In general, $Y$ and $S$ will be correlated at the unit level outside of any particular experiment (Cunningham and Kim, 2020). For example, highly engaged users ($S$) are also likely to be more retentive ($Y$), regardless of whether interventions that increase engagement also increase retention. Thus, we distinguish between the true associations between the underlying treatment effects and the observed associations between treatment effects estimated in finite experiments (Bibaut et al., 2024). This finite-sample noise amplifies the winner’s curse for proxy metrics and motivates the use of our CV estimator.
@@
(7) $R_{iD_{S}(\mathbf{O}^{i})}=\begin{cases}\tau_{Y}^{i}&\text{if $\hat{\tau}_{S}^{i}>0$}\\ 0&\text{otherwise}.\end{cases}$
@@
4.2. Closed-Form Analysis
@@
(8) $(\tau_{R}^{i},\tau_{S}^{i})\sim\mathcal{N}\left(\mathbf{0},\Lambda\right).$
@@
(9) $\mathbb{E}[R_{iD_{S}}]\propto\frac{\Lambda_{12}}{\sqrt{\Lambda_{22}+2\Omega_{22}/M}},$
which intuitively is increasing in the covariance in true treatment effects. [^5] See Appendix 7.1 for derivations of Equations 9-11.
@@
(10) $\mathbb{E}[\hat{R}^{naive}_{iD_{S}}]\propto\frac{\Lambda_{12}+2\Omega_{12}/M}{\sqrt{\Lambda_{22}+2\Omega_{22}/M}}.$
Therefore, the naive estimator can prefer proxy metrics whose correlation with the north star is driven more by noise than by true treatment effects. A standard example is clickbait: Although highly engaged users may be more retentive outside of any particular experiment ($\Omega_{12}$ large), interventions that merely increase clickbait will not have the desired effect on retention ($\Lambda_{12}$ small).
Note that the risk of bias is greater when the scale of true treatment effects is very small relative to the unit-level sampling variance, as is often the case in digital experiments (Larsen et al., 2024). Thus, the naive estimator will skew towards sensitive “upper funnel” metrics such as page visits, as compared to “lower funnel” metrics that are harder to move (e.g., signups or checkouts) but more strongly correlated with north star business metrics like revenue.
@@
(11) $\mathbb{E}[\hat{R}^{CV}_{iD_{S}}]\propto\frac{\Lambda_{12}}{\sqrt{\Lambda_{22}+2\Omega_{22}/M_{p}}},$
where $M_{p}=M(P-1)/P$. Like the true reward, this expectation is also insensitive to the noise covariance $\Omega_{12}$, depending only on the covariance between the true treatment effects. It is also negatively biased due to the division in (11) by $M_{p}<M$, although this can be mitigated by increasing $P$.
Figure 1. Expected Values of True Reward and Estimators. While the true expected reward only depends on the covariance of true treatment effects, the naive estimator is sloped and can prefer “Bad” proxies to “Good” proxies. The CV estimator is negatively biased (diluted) but still ranks proxies correctly.
Figure 1 illustrates these dynamics. In the top panel, we plot the true reward $\mathbb{E}[R_{iD_{S}}]$ as the correlation between true treatment effects ($y$-axis) and measurement error ($x$-axis) varies from -1 to 1.666The other parameter values are set to the same values as in our simulation below. Regions of the same shade represent level sets (the same value of the reward). As the panel shows, the true expected reward is only sensitive to the correlation in true treatment effects.
@@
Lastly, the bottom panel of Figure 1 plots the expected value of our CV estimator with just two folds. The flat level sets demonstrate that the CV estimator will rank correctly based on the true reward, although their shade is diluted due to the secondary source of bias in the CV-estimated reward. In practice, this bias is easily mitigated by increasing the number of folds.
4.3. Simulation Study
@@
As can be seen from Equations 9 and 10, the relative bias of the naive estimator is proportional to the ratio of the covariance in measurement error to the covariance in true treatment effects $\Omega_{12}/\Lambda_{12}$. Letting $\sigma_{S}$ denote the standard deviation of $S$ and $\sigma_{\tau_{S}}$ the standard deviation of $\tau_{S}$, the relative bias is therefore proportional to $\sigma_{S}/\sigma_{\tau_{S}}$, the noise-to-signal ratio of $S$. Therefore, as treatment effects on the proxy metric become infinitely small relative to their measurement error, the relative bias of the naive estimator also tends to infinity. This is illustrated in Figure 2, where we simulate the bias of the naive estimator as the scale of the noise $\sigma_{S}$ tends to infinity with $\sigma_{\tau_{S}}$ fixed. As the plot shows, the relative bias of the naive estimator is also unbounded. In contrast, because the CV estimator is biased towards zero, its relative bias is negative for positive rewards and bounded between $-100\%$ and $100\%$.
Figure 2. Bias of Naive and CV Estimators vs. Noise-to-Signal Ratio of $S$. As the scale of the measurement error in $S$ tends to infinity for a fixed scale of treatment effects on $S$, the bias of the naive estimator tends to infinity, whereas the relative bias of the CV estimator is bounded.
@@
Figure 3. Asymptotics as $M\to\infty$. The light-opacity dashed lines denote the truth, while the solid lines denote estimators. While the bias of both estimators goes to zero as the number of units per treatment arm $M\to\infty$, the naive estimator is much more biased at small values of $M$, which can lead to the wrong choice of proxy.
@@
Figure 4. Asymptotics as $N\to\infty$. The light-opacity dashed lines denote the truth, while the solid lines denote estimators. Increasing the number of experiments $N$ without increasing the size of each experiment only increases the bias of the naive estimator.
5. Real-World Application
In this section, we describe the application of our method to successfully advocate for a new experimentation decision rule at Netflix. This work is part of a larger project to develop proxy metrics using meta-analysis, initially described in (Bibaut et al., 2024). In that paper, we developed methods for fitting linear structural models of treatment effects using data from past experiments. We deployed these methods to construct an improved proxy metric as a linear combination of component metrics, which include the previous status quo metric. We then used techniques in this paper to demonstrate the favorability of the resulting metric relative to the previous decision metric. Our method was essential to convincing our stakeholders to adopt the new decision rule. This led directly to the adoption and implementation of the new proxy metric on Netflix’s experimentation platform, where it has been computed for 257 A/B tests as of this writing.
5.1. Setup
The data for this case study come from 123 A/B past tests at Netflix. We chose these tests to be representative of the primary testing area that we wanted our new proxy metric to inform. Specifically, we sampled tests of in-app interventions aimed at improving engagement, such as innovations to our recommendation and personalization algorithms.
Letting $Y$, $Z_{0}$, and $Z_{1}$ denote the north star metric, status quo proxy metric, and the new “challenger” proxy metric, respectively, we estimated the cumulative returns to a handful of candidate launch rules: [^9] We also experimented with a number of other combinations (such as different tie-breakers and requiring statistical significance in both $Z_{0}$ and $Z_{1}$), which we omit for brevity.
  1. (1)
@@ 5.2. Results
Figure 5. Estimated Cumulative Returns to Decision Rules Across 123 A/B Tests at Netflix. We applied the methods in this paper to estimate the naive and cross-validated cumulative returns to three different decision rules at Netflix, contrasting two challenger rules to the status quo rule. We find that a new proxy metric would have increased cumulative returns by over $33\%$ ($p$-values between $0.07$ and $0.09$).
@@
Figure 5 shows our results by plotting the naive (“No Folds”) estimate of the cumulative returns (red crosses) alongside our CV estimates under increasingly numerous folds (blue circles). Error bars represent 95% confidence intervals. To maintain confidentiality, we scale all values by baseline gains (the “No Folds” variant for Rule 1).
@@
As expected, the naive estimator systematically overestimates the cumulative returns of past experiments due to the winner’s curse. Although cross-validation could overshrink these estimates, the estimates remain stable as we increase the number of folds, indicating that the negative bias of the CV estimator is small in our application.
@@
As the figure shows, our methods can also influence qualitative comparisons between decision rules. The naive ranking places Rule 3 far above Rule 2, while the CV estimator makes the two rules appear comparable. This is likely due to a combination of $Z_{0}$ having a more severe correlated errors problem than $Z_{1}$, as well as the slightly higher Type I error rate incurred by using an either-or launch rule. Altogether, the results demonstrate that the choice of estimator can meaningfully alter both quantitative (estimates of cumulative impact) and qualitative (choice of rule) business conclusions.
@@
We find that both “challenger” rules would meaningfully outperform the status quo. In particular, we estimate that adopting Rule 3 would have increased north star metric gains in excess of 33%, with $p$-values between 0.07 and 0.09 depending on the number of folds. Upon closer inspection, we found that these improvements stemmed largely from a few high-impact tests in which our new proxy metric disagreed sharply with the previous decision metric. Individually, these tests were underpowered for the north star metric, but collectively they accounted for a substantial improvement. Although marginally significant under traditional frequentist thresholds, the large gains at stake and modest implementation costs gave us confidence to recommend Rule 3. These estimated gains, as well as the explainability of our methodology, led directly to the adoption of this new launch rule by decision makers. Our rule has since become recognized as the best practice for resolving metric disagreements in experiments.
6. Conclusion
@@
This paper proposes to evaluate candidate decision rules using their cumulative returns to business north star metrics in past experiments. In our experience, cumulative returns are an intuitive evaluation metric that resonates with less technical stakeholders, making them a compelling quantity to report across a broad set of applications. Still, accurately estimating cumulative returns can be challenging when treatment effects are small relative to measurement error, as is often the case in digital experimentation, including at Netflix (Bibaut et al., 2024; Larsen et al., 2024; Ejdemyr et al., 2024).
@@
We demonstrate theoretically and empirically that cross-validation can effectively estimate the reward of a decision rule in such conditions. In contrast, naive approaches that reuse the same data for both decision rule selection and evaluation lead to classical winner’s curse biases, even when experimenters are optimizing for proxy metrics rather than north star metrics.
@@
An implicit assumption underlying this work is the independence of treatment effects across tests: The difference in rewards between two treatment arms within an experiment cannot depend on the launch decisions of other experiments. This assumption could be violated if experiments implement redundant interventions or if previous launch decisions radically affect the design and impact of future tests. We believe that this is a reasonable approximation in our setting, which is characterized by many modest interventions across diverse domains and experiments that are typically planned far in advance. That being said, one valuable direction for future research is to explore estimators that allow for dependence across experiments, for example by encoding dependencies between tests in cross-validation. Relatedly, we also implicitly assume a degree of stationarity, such that rules optimized for past returns will also be useful for future experiments. This assumption can be monitored by re-running the analysis periodically over time.
@@
A second, more applied direction in which to extend this research is to evaluate other dimensions of experiment decision-making beyond the choice of outcome metric. For example, recent papers have debated whether statistical significance thresholds should be more restrictive or permissive (Kohavi and Chen, 2024; Sudijono et al., 2024). The methodology outlined here provides a principled way for firms to evaluate the returns to different thresholds using actual experiments, and can easily be extended to include costs.
@@
Lastly, another direction for future research is to learn good decision rules (rather than pick the best rule from an ex ante-defined set). Here, an interesting trade-off is between optimality and interpretability: a good decision rule should increase the returns to experimentation, but also be coherent enough to be trusted, understood, and remembered by decision-makers with varying statistical aptitudes. We leave this to future work.